OpenStudy (anonymous):

Hello. Is it possible to cancel out terms in derivatives? For example, (dX/dY)/(dX/dZ). Please, note that X reacts in the same way to the changes in Y and Z. In this case, could we write it like this: (dX/dY)/(dX/dZ)=dZ/dY ? thank you in advance.

OpenStudy (anonymous):

yes, it is possible. When working with parametric equations you can take \[\frac{ \frac{ dy }{ dt } }{ \frac{ dx }{ dt } } = \frac{ dy }{ dx }\]

OpenStudy (anonymous):

thanx a lot